OpenStudy (anonymous):

Write an equation that represents the problem, then solve the problem. Seven-fourths decreased by a number is 2. Find the number

OpenStudy (mathlegend):

@kbevel So we know.... 7/4 decreased by a name: -x is 2: = 2

OpenStudy (mathlegend):

@kbevel can you put that together?

OpenStudy (anonymous):

no

OpenStudy (mathlegend):

7/4 - x = 2

OpenStudy (mathlegend):

@kbevel do you understand how I did this so far?

OpenStudy (mathlegend):

@kbevel ???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sorry everything froze on me

OpenStudy (anonymous):

ok do I need to make the 2 a fraction and such?

OpenStudy (mathlegend):

7/4 - x = 2 can you solve this?

OpenStudy (anonymous):

subtract 7/4 from both sides?

OpenStudy (mathlegend):

Yes

OpenStudy (anonymous):

ok x=1/4

OpenStudy (mathlegend):

Yes.

OpenStudy (anonymous):

ok ty